OpenStudy (anonymous):

Find the intersections of y=x+2 and y^2 =4x

OpenStudy (dumbcow):

solve 2nd equation for y --> y = +-sqrt(4x) = +-2sqrtx set 2 equations equal, solve forx +-2sqrtx = x+2 4x= (x+2)^2 4x = x^2 +4x +4 0 = x^2 +4 -4 = x^2 can't take sqrt of neg number therefore, No solution, these functions never intersect

OpenStudy (anonymous):

WOW I HATE IT WHEN THERE IS NO SOLUTION THANKS!!!!